# Observation of Fermi acceleration with cold atoms

### v1

@@ Abstract
Cosmic rays are deemed to be generated by a process known as “Fermi acceleration”, in which charged particles scatter against magnetic fluctuations in astrophysical plasmas. The process itself is however universal, has both classical and quantum formulations, and is at the basis of dynamical systems with interesting mathematical properties, such as the celebrated Fermi-Ulam model. Despite its effectiveness in accelerating particles, Fermi acceleration has so far eluded unambiguous verifications in laboratory settings. Here, we realize the first fully controllable Fermi accelerator by colliding ultracold atoms against engineered movable potential barriers. We demonstrate that our Fermi accelerator, which is only 100 $\mu$m in size, can produce ultracold atomic jets with velocities above half a meter per second. Adding dissipation, we also experimentally test Bell’s general argument for the ensuing energy spectra, which is at the basis of any model of cosmic ray acceleration. On the one hand, our work effectively opens the window to the study of high energy astrophysics with cold atoms, offering new capabilities for the understanding of phenomena such as diffusive acceleration at collisionless shocks. On the other, the performance of our Fermi accelerator is competitive with those of best-in-class accelerating methods used in quantum technology and quantum colliders, but with substantially simpler implementation and virtually no upper limit.
@@
The cornerstone mechanism for the origin of cosmic rays was proposed by E. Fermi [3, 4]: when a particle undergoes a head-on (tail-on) elastic collision with a moving magnetic mirror it gains (loses) kinetic energy, but since head-on encounters are more likely than tail-on ones (think about inverse-Compton scattering), on average the particle’s energy must increase with time. While in astrophysical plasmas this Fermi acceleration is realized by scattering charged particles against magnetic fluctuations, the process itself is relevant for a wide range of terrestrial classical [5, 6, 7] and quantum [8, 9, 10] systems, and has inspired several mathematical models with unique properties [11, 12, 13]. In the past two decades, kinetic simulations of astrophysical plasmas exploiting the growing power of modern supercomputers for solving Maxwell’s equations coupled with a phase-space description of particles, have been able to shed light on how acceleration may occur in astrophysical shocks [14, 15, 16, 17, 18], in magnetic reconnection [19, 20, 21], and in turbulence [22, 23, 24]. Experimentally, several impressive endeavors are working towards producing particle acceleration at shock in laboratory settings [25, 26, 27, 28, 29]. However, to date only hints of deviations from thermal distributions have been reported. In fact, sustaining a plasma laser experiment for many acceleration cycles to detect extended spectra of accelerated particles is still out of reach.
@@
Our Fermi accelerator is implemented with ultracold 87Rb atoms and potential barriers generated by laser light. The barriers are dynamically shaped with a digital micromirror device [30], and their height is controlled with the power and detuning of the laser beam [30]. The atoms are initially trapped in a crossed beam dipole trap, which is deformed into a one dimensional waveguide along the direction of the accelerator before the acceleration process begins. This is done to support the atomic jet against gravity [30]. In Fig. 2a) we report a typical experimental sequence for an accelerator with $A/k_{B}=0.7A_{0}/k_{B}\simeq 170$ $\mu$K, where $k_{B}$ is the Boltzmann constant, observed with absorption imaging. The moving barrier, whose position is approximately indicated by the black dashed line, hits the ultracold sample that is initially at rest. The center of mass of the atomic cloud then moves with velocity $2v_{i}$ towards the static barrier, indicated by the yellow dashed line. Similarly to Fig. 1b), the subsequent microscopic dynamics cannot be resolved in situ, due to the spread of the cloud in momentum and real space. Nonetheless, the net effect of the Fermi acceleration acting on all the atoms manifests itself with the atoms being ejected from the accelerator. As a result of the acceleration process, we observe atomic jets traveling at high speed in the waveguide, as shown in Fig. 1c).
@@
We finally utilize our setup to experimentally test the famous Bell’s argument [34], which is one of the cornerstones of modern high-energy astrophysics, where extended power-law distributions are routinely observed, both in the cosmic radiation (electrons, nuclei, and neutrinos) and in the non-thermal emission from astrophysical objects. Bell argued that power-law distributions generally arise as the result of a combination of energy gain and particle escape from the accelerator. In our system, Bell’s argument predicts that the energy spectrum inside the accelerator must have the form $N_{n}=N_{1}(E_{n}/E_{1})^{-Q}$ with $Q=-\ln(P)/\ln(G)$, where $N_{1}(E_{1})$ and $N_{n}(E_{n})$ are the number of atoms (average energy) after one and $n$ collisions with the moving barrier respectively. $P$ and $G$ are the survival probability and energy gain after each collision [30]. To verify this prediction, we slightly modified the experimental sequence so to measure the number of atoms left and their average velocity after each collision with the barriers [30]. By controlling the detuning and power of the laser light that generates the potential barriers, we are able to tune the light-atom scattering rate $\Gamma_{s}$, which in turn determines $P$, leaving all the other parameters of the experiment unchanged. In Fig. 4a) we report our results on the number of atoms left in the accelerators as a function of their average energy, for different values of $\Gamma_{s}$. Our data are well fitted by the power-law function described above (solid lines), with $Q=-\Theta/\ln(G)$, where $\Theta$ is the fitting parameter and $G$ is a function of $E_{n}/E_{1}$ [30]. In general, atom losses in our system are a combination of losses coming from the light-atom scattering discussed above and other loss mechanisms that don’t depend on the detuning, the most relevant being three-body collisions. For this reason, in our experiment we expect the survival probability to have the form $P=(\gamma_{0}+\xi\Gamma_{s})^{-1}$, where $\gamma_{0}$ accounts for all the detuning-independent loss mechanisms and $\xi$ for the effective interaction time of the atoms with the barriers. As shown in Fig. 4b), the scaling of the $\Theta$ parameter observed in our experiment is remarkably compatible with the functional dependence $\Theta=\ln(\gamma_{0}+\xi\Gamma_{s})\equiv\ln(P)$ (solid line), as predicted by Bell’s argument.
@@
Our results show that cold atom technology is an ideal platform to implement and study Fermi acceleration processes and related phenomena. For example it can be used as a testbed for classical, dissipative and quantum Fermi-Ulam models [11, 12, 13, 8], and can provide new opportunities towards the understanding of complex mechanisms such as diffusive shock acceleration in astrophysical plasmas [35, 36, 37, 34, 38]. In this direction, particularly appealing is the ability of cold atom experiments to engineer the interactions between the atoms [39, 40]. Our methods can for example be straightforwardly applied to ultracold neutral plasmas [41, 42], where some of the conditions found in astrophysical settings could be obtained [43]. In the context of quantum technology, our Fermi accelerator enables to obtain samples with large $E_{k}$ and low effective temperatures, that are desirable features in quantum chemistry [44, 45, 46] and quantum colliders applications [47, 48, 49]. With respect to other record-achieving techniques in atom optics [50], a Fermi accelerator is characterized by a remarkably simple implementation, which makes it ideally suited for in-trap atomtronics applications [51]. Fermi acceleration can be also used to direct and split atomic samples into two counter-propagating clouds [30], thus providing complementary features to existing techniques [31, 32]. Finally, the study of quantum Fermi acceleration [8, 30] could be a promising avenue for quantum information science, as there is concrete potential for the development of new tools for the manipulation of quantum wavepackets .

### v2

@@ Abstract
Cosmic rays are deemed to be generated by a process known as Fermi acceleration, in which charged particles scatter against magnetic fluctuations in astrophysical plasmas. The process itself is however universal, has both classical and quantum formulations, and is at the basis of dynamical systems with interesting mathematical properties, such as the celebrated Fermi-Ulam model. Despite its effectiveness in accelerating particles, Fermi acceleration has so far eluded unambiguous verifications in laboratory settings. Here, we realize a fully controllable Fermi accelerator by colliding ultracold atoms against engineered movable potential barriers. We demonstrate that our Fermi accelerator, which is only 100 $\mu$m in size, can produce ultracold atomic jets with velocities above half a meter per second. Adding dissipation, we also experimentally test Bell’s general argument for the ensuing energy spectra, which is at the basis of any model of cosmic ray acceleration. On the one hand, our work effectively opens the window to the use of cold atoms to study phenomena relevant for high energy astrophysics. On the other, the performance of our Fermi accelerator is competitive with those of best-in-class accelerating methods used in quantum technology and quantum colliders, but with substantially simpler implementation and virtually no upper limit.
@@
The cornerstone mechanism for the origin of cosmic rays was proposed by E. Fermi [3, 4]: when a particle undergoes a head-on (tail-on) elastic collision with a moving magnetic mirror it gains (loses) kinetic energy, but since head-on encounters are more likely than tail-on ones (think about inverse-Compton scattering), on average the particle’s energy must increase with time. While in astrophysical plasmas this Fermi acceleration is realized by scattering charged particles against magnetic fluctuations, the process itself is relevant for a wide range of terrestrial classical [5, 6, 7] and quantum [8, 9, 10] systems, and has inspired several mathematical models with unique properties [11, 12, 13]. In the past two decades, kinetic simulations of astrophysical plasmas exploiting the growing power of modern supercomputers for solving Maxwell’s equations coupled with a phase-space description of particles, have been able to shed light on how acceleration may occur in astrophysical shocks [14, 15, 16, 17, 18], in magnetic reconnection [19, 20, 21, 22], and in turbulence [23, 24, 25]. Experimentally, several impressive endeavors are working towards producing particle acceleration at shock in laboratory settings [26, 27, 28, 29, 30]. However, to date only hints of deviations from thermal distributions have been reported. In fact, sustaining a plasma laser experiment for many acceleration cycles to detect extended spectra of accelerated particles is still out of reach.
@@
Our Fermi accelerator is implemented with ultracold 87Rb atoms and potential barriers generated by laser light. This latter indeed produces an optical dipole force [32] whose strength can be controlled with the power and detuning of the laser beam [31]. The barriers are dynamically generated by using a digital micromirror device that shapes the profile of the laser beam [31]. The atoms are initially trapped in a crossed beam dipole trap, which is deformed into a one dimensional waveguide along the direction of the accelerator before the acceleration process begins. This is done to support the atomic jet against gravity [31]. In Fig. 2a) we report a typical experimental sequence for an accelerator with $A/k_{B}=0.7A_{0}/k_{B}\simeq 170$ $\mu$K, where $k_{B}$ is the Boltzmann constant, observed with absorption imaging. The moving barrier, whose position is approximately indicated by the black dashed line, hits the ultracold sample that is initially at rest. The center of mass of the atomic cloud then moves with velocity $2v_{i}$ towards the static barrier, indicated by the yellow dashed line. Similarly to Fig. 1b), the subsequent microscopic dynamics cannot be resolved in situ, due to the spread of the cloud in momentum and real space. Nonetheless, the net effect of the Fermi acceleration acting on all the atoms manifests itself with the atoms being ejected from the accelerator. As a result of the acceleration process, we observe atomic jets traveling at high speed in the waveguide, as shown in Fig. 2b). Note that the direction of the atomic jet can be controlled with the $A/A_{0}$ ratio. For $A>A_{0}$, the atoms escape in the opposite direction (upstream) [31].
@@
We finally utilize our setup to experimentally test the famous Bell’s argument [36], which is one of the cornerstones of modern high-energy astrophysics, where extended power-law distributions are routinely observed, both in the cosmic radiation (electrons, nuclei, and neutrinos) and in the non-thermal emission from astrophysical objects. Bell argued that power-law distributions generally arise as the result of a combination of energy gain and particle escape from the accelerator. In our system, Bell’s argument predicts that the energy spectrum inside the accelerator must have the form $N_{n}=N_{1}(E_{n}/E_{1})^{-Q}$ with $Q=-\ln(P)/\ln(G)$, where $N_{1}(E_{1})$ and $N_{n}(E_{n})$ are the number of atoms (average energy) after one and $n$ collisions with the moving barrier respectively. $P$ and $G$ are the survival probability and energy gain after each collision [31]. To verify this prediction, we slightly modified the experimental sequence so to measure the number of atoms left and their average velocity after each collision with the barriers [31]. By controlling the detuning and power of the laser light that generates the potential barriers, we are able to tune the light-atom scattering rate $\Gamma_{s}$, which in turn determines $P$, leaving all the other parameters of the experiment unchanged. For this set of data we work close to resonance to increase $\Gamma_{s}$ [31]. In Fig. 4a) we report our results on the number of atoms left in the accelerators as a function of their average energy, for different values of $\Gamma_{s}$. Our data are well fitted by the power-law function described above (solid lines), with $Q=\Theta/\ln(G)$, where $\Theta$ is the fitting parameter and $G$ is a function of $E_{n}/E_{1}$ [31]. In general, atom losses in our system are a combination of losses coming from the light-atom scattering discussed above and other loss mechanisms that don’t depend on the detuning, the most relevant being three-body collisions. For this reason, in our experiment we expect the survival probability to have the form $P=(\gamma_{0}+\xi\Gamma_{s})^{-1}$, where $\gamma_{0}$ accounts for all the detuning-independent loss mechanisms and $\xi$ for the effective interaction time of the atoms with the barriers. As shown in Fig. 4b), the scaling of the $\Theta$ parameter observed in our experiment is remarkably compatible with the functional dependence $\Theta=\ln(\gamma_{0}+\xi\Gamma_{s})\equiv-\ln(P)$ (solid line), as predicted by Bell’s argument.
Our results show that cold atom technology is an ideal platform to implement and study Fermi acceleration processes and related phenomena. For example it can be used as a testbed for classical, dissipative and quantum Fermi-Ulam models [11, 12, 13, 8]. While several astrophysical phenomena will probably remain out of reach for table-top experiments, by combining the many tools available in cold atom systems one could investigate some of the open questions concerning particle acceleration at shock [37, 38, 39, 36, 40], but also emerging phenomenology linked to propagating Alfvén waves [41], for example using multiple moving barriers. Our methods could be straightforwardly applied to ultracold neutral plasmas [42, 43], where some of the conditions found in astrophysical settings could be obtained [44]. Particularly interesting is the possibility of controlling the strength and the nature of the interactions [45], which is a major challenge in many-body numerical simulations. The experiment could test how different kinds of interactions affect the acceleration rate and therefore the maximum energy attainable. In addition, one could compare diffusion and Levy flights regimes, study super-diffusion, and implement variations or periodicity in the speed of the scattering centers, which are all important aspects in astrophysical environments. The complete control over the potential barriers is another key element that could be exploited: for example one could dynamically change the height and the speed of the barriers to accelerate in different ways different classes of atoms and simulate the effect of a cosmic ray shock precursor. One possible extension of our experimental setup is the study of stochastic (or second-order) Fermi acceleration, which is potentially important in astrophysics [46, 47]. Differently from first-order acceleration in shocks, stochastic acceleration does not lead to universal power laws in energy, and the theory that connects the statistics of scattering agents and particle energization is an active area of research [48]. Using a superposition of potentials with a prescribed power spectrum and measuring the ensuing particle spectrum may inform models of cosmic ray acceleration in both galactic and extragalactic environments.
@@
In the context of quantum technology, our Fermi accelerator enables to obtain samples with large $E_{k}$ and low effective temperatures, that are desirable features in quantum chemistry [49, 50, 51] and quantum colliders applications [52, 53, 54]. With respect to other record-achieving techniques in atom optics [55], a Fermi accelerator is characterized by a remarkably simple implementation, which makes it ideally suited for in-trap atomtronics applications [56]. Fermi acceleration can be also used to direct and split atomic samples into two counter-propagating clouds [31], thus providing complementary features to existing techniques [33, 34]. Finally, the study of quantum Fermi acceleration [8, 31] could be a promising avenue for quantum information science, as there is concrete potential for the development of new tools for the manipulation of quantum wavepackets .